# SLIM-Brain: A Data- and Training-Efficient Foundation Model for fMRI Data Analysis

### v1

@@ 1 Introduction
More recently, there has been a growing interest in developing deep learning based foundation models for fMRI data analysis, inspired by their remarkable performance in Computer Vision and Natural Language Processing tasks. The idea is to pre-train (usually in a self-supervised way) a large model to learn general-purpose representations of brain activity on vast neuroimaging data, which can then be adapted to diverse downstream tasks with limited labeled data (e.g., via fine-tuning). Broadly speaking, current efforts on building foundation models, or more generally applying deep learning techniques in fMRI data analysis can be categorized into two threads: atlas-based and atlas-free approaches. The former continues the traditional paradigm of summarizing voxel-level signals into predefined brain regions (caro2023brainlm; assran2023self; yang2024brainmass), therefore leveraging anatomical priors and improving interpretability. However those methods bear several key limitations: i) there is no universally optimal atlas: performance on different downstream tasks may highly depend on parcellation choices, where results across studies using heterogeneous atlas are not directly comparable (wang2025dca; salehi2020there); ii) averaging based on any pre-defined atlas will inevitably discard important voxel-level information - as a result such models often need to be trained on very large cohorts (e.g., $\sim 60k$ sessions) to perform well (dong2024brain; caro2023brainlm) (see Fig. 1); and iii) any analysis using such models will be confined to the parcel resolution of the atlas used, where probing within-region structures (e.g., isolating amygdalar subnuclei during fear conditioning) is impossible (Wen2022TemporallyAA).
@@
On the other hand, atlas-free methods (zhao2018modeling; vu2020fmri; nguyen2020attend) aim to learn directly from voxel-level data without imposing ROI boundaries of certain atlas, allowing the learned models to capture fine-grained functional patterns and potentially discover novel brain organization. However, existing atlas-free approaches have typically been developed for specific tasks rather than as general-purpose foundation models, due to their prohibitively high training cost. The deep learning architectures underlying those models, such as the widely adopted Vision Transformers (ViT), incur quadratic cost (both memory and compute) with respect to input dimensions. This makes it nearly impossible to pre-train them on large-scale fMRI data, and thus building atlas-free foundation models still remains an open challenge. Recent work has explored efficient variants such as Shifted-window (Swin) Transformer (kim2023swift; Sun2025VoxelLevelBS; Kwon2024PredictingTB; peng2025whole), but at each timestamp they still feed the entire dense fMRI volume into the encoder, wasting resources on those voxels ($\approx$70%) outside the brain with no valid signal. Similar inefficiencies also arise along the temporal axis, where existing approaches either only train with data pertain to specific task states, e.g., extracting $\sim$30 timestamps (shi2023self; Sun2025VoxelLevelBS; rosenman2024pre), or consider sliding windows with limited sizes, where data within each window is processed independently through the model, with results aggregated thereafter (kim2023swift)(e.g., as shown in Fig. 2b). In practice neither is ideal: the former obviously hurts generalization capabilities of the model beyond task states, while the latter evenly processes every window - may lose focus on those truly important data segments - leading to inferior performance comparing to some of the recent atlas-based approaches in out-of-distribution cases (dong2024brain).
@@ 2 Related Work
Foundation Models for fMRI Analysis. fMRI signals reflect ongoing brain states and cognitive processes. Early work framed decoding as supervised classification or regression on activity patterns, which often produced task-specific features with limited out-of-distribution generalization (song2008brain; horikawa2017generic; ye2023explainable; zhao2018modeling; vu2020fmri; nguyen2020attend). More recently, the field has pivoted to task-agnostic foundation models trained with self-supervised pipelines on unlabeled data, aiming for representations that transfer across tasks and datasets. One prominent line of work pools voxel signals within a predefined atlas to obtain region-wise time series and then learns objectives on those ROI tokens—either generative (masked reconstruction, as in BrainLM) or latent prediction (as in Brain-JEPA) (caro2023brainlm; dong2024brain). These approaches are memory-efficient and scalable, but their quality can depend on the chosen parcellation and its granularity. A related direction converts the atlas-reduced time series into pairwise or higher-order functional graphs and optimizes unsupervised objectives on these structures (yang2024brainmass; thapaliya2024dsamad; han2025hypergraph). For example, BrainMass augments networks by randomly dropping time points from the BOLD signal during training to encourage robustness (yang2024brainmass), and hypergraph formulations have been explored to capture higher-order relationships (han2025hypergraph). Complementing these atlas-based routes, volumetric encoders operate directly on 4-D fMRI volumes, thereby avoiding atlas-induced bias and preserving fine spatial detail (Peng2025WholebrainTR; Sun2025VoxelLevelBS; Kwon2024PredictingTB; kim2023swift).In practice, computational and memory constraints often necessitate training on short, task-aligned excerpts rather than full-length recordings. Typical applications include task classification (shi2023self), brain-state decoding (Sun2025VoxelLevelBS), and stress prediction (rosenman2024pre). Turning these systems into a general-purpose foundation model remains a major challenge.
@@
Efficient Vision Transformers. Vanilla Vision Transformers (ViTs) apply global self-attention over all tokens. The cost scales quadratically with token count and the model offers weak spatial inductive bias (dosovitskiy2020image). This is impractical for high-resolution or long 4-D fMRI sequences, where a single scan can yield millions of voxel tokens across time. Hierarchical Transformers address these issues by progressively downsampling tokens and widening channels (e.g., MViT, Swin) (fan2021multiscale; liu2021swin). For example, Swin restricts attention to local windows and cyclically shifts the partition between layers, which stabilizes memory and accuracy on dense inputs. However, it limits global context to multi-layer message passing, ties the model to a dense, regular lattice, and introduces engineering overhead that complicates masked-token pretraining. These drawbacks are magnified in fMRI: spatial grids are high resolution and a large fraction of voxels are non-brain background. Hiera shows that many hand-crafted components are unnecessary. A minimal hierarchical ViT by pretext task achieves superior speed/accuracy trade-offs and is naturally compatible with MAE-style pretraining (ryali2023hiera). For voxel-level 4-D fMRI, this is particularly attractive: with MAE-style masking, non-informative background can be excluded from the encoder, reducing GPU memory and improving throughput while preserving fine-grained brain signals.
@@ 4.1 Dataset
We leverage self-supervised pretraining on four public neuroimaging datasets: a single HCP session (van2013wu), CHCP (ge2023increasing), and two releases from the Amsterdam Open MRI Collection (AOMIC)—PIOP1 and PIOP2 (snoek2021amsterdam); the Adolescent Brain Cognitive Development (ABCD) Study casey2018adolescent. We use 70% of the data (4129 sessions) for training, with the remainder reserved for validation (10%) and testing (20%). External validation spans four datasets: the Autism Brain Imaging Data Exchange (ABIDE) (di2014autism); the ADHD-200 Sample (ADHD) (adhd2012adhd); Alzheimer’s Disease Neuroimaging Initiative (ADNI) jack2008alzheimer; Parkinson’s Progression Markers Initiative (PPMI) marek2011parkinson. See Appendix  A.1 and  A.2 for preprocessing and dataset details.
@@ 4.2 Fine-tuning results
We fine-tune our model and evaluate it on out-of-distribution datasets, as summarized in Table 1 and Appendix A.3. We benchmark against seven representative fMRI models—BrainNetCNN (kawahara2017brainnetcnn), BrainGNN li2021braingnn, BrainLM (caro2023brainlm), BrainMass (yang2024brainmass), Brain-JEPA (dong2024brain), NeuroSTORM li2025towards and SwiFT kim2023swift—using the authors’ released codes or checkpoints (pretrained on 1010k–6464k samples). Note that BrainMass includes these datasets in its pre-training corpus, whereas our method utilizes them strictly for fine-tuning. We adopted the medium-sized model (45M parameters) trained on 4,129 sessions as the default configuration for the experiments. Across all the OOD evaluations, our model consistently outperforms the baselines, indicating that it learns domain-invariant representations that transfer to unseen datasets and tasks with minimal adaptation. Task description (A.2), statistical method (A.3), internal experiment results (A.4), linear probing results results (A.5), configuration details for ours (A.6) and baselines (A.7) are provided in the Appendix.
@@ 4.3 Ablation Study
To avoid ingesting full-length fMRI sequences, we use the global MAE to obtain a coarse descriptor and to score temporal windows; only the top-ranked windows are fed to the 4D encoder (Table 3). We use five consecutive frames to form one window, and multiple such windows are grouped into a set as the input to the voxel-level model. With a per-frame storage layout, we can load arbitrary time indices on demand, cutting $\sim 80\%$ of 4D data I/O and compute when selecting the top $20\%$ of windows, comparing sliding windows average method (kim2023swift). We compare Top-$k$ selection (mutual masked reconstruction) against Random selection under different frame budgets on HCP sex classification with 1K-size pre-trainging model. Top-$k$ consistently achieves higher accuracy and F1 without full-sequence 4D fMRI reads.
@@
We assess the impact of our design decisions on memory and accuracy by 1K-size pretraining model with HCP dataset in Table. 4. First, employing a 4D Hiera encoder with unit-wise attention excludes background and masked units from the forward pass, substantially reducing the memory footprint. Because fMRI signals are not directly interpretable like natural images, we prioritize representation quality over pixel-space fidelity and adopt a JEPA objective, dispensing with a heavy reconstruction decoder. With 200-frame inputs and Top-$k$ selection, our design reduces peak GPU memory from $\sim$8 GB to $\sim$2.4 GB while improving accuracy. Meanwhile, although Hiera–MAE and Hiera–JEPA have comparable per-sample GPU memory footprints, the JEPA variant achieves substantially higher throughput. By contrast, kim2023swift processes ten 20-frame windows sequentially and averages their predictions; each 20-frame chunk consumes $\approx 3.2\,\mathrm{GB}$ of GPU memory, so handling all windows inflates the per-sample memory footprint. Likewise, Kwon2024PredictingTB employ a Swin-UNet–based model for task prediction; although it is not a foundation model and does not reconstruct to the raw signal space, it still requires $\sim 40\,\mathrm{GB}$ of GPU memory to process 30 volumes with a batch size of four. Details are in Appendix A.10
@@ 4.4 Model interpretation
We examine whether SLIM-Brain’s voxel-level representations are neurobiologically meaningful. First, we use Neurosynth—a large-scale, automated meta-analytic platform that aggregates findings from thousands of fMRI studies (yarkoni2011large)—to obtain disease-associated meta-analytic maps (e.g., ADHD). Next, we derive fine-grained attention maps from the 4D Hiera encoder by projecting token-level attention back to voxel space. (Implementation details are provided in the Appendix A.11.) Integrated-gradients–derived key regions from SLIM-Brain substantially overlap with Neurosynth meta-analytic distributions (Fig. 5). The highlighted regions include dorsolateral prefrontal cortex and anterior cingulate (executive control), inferior parietal and precuneus/posterior cingulate (default mode network), as well as striatal areas, consistent with ADHD-related fronto-striatal and fronto-parietal dysfunctions. These observations indicate that the learned features are not only predictive but also align with established neurobiological patterns, offering interpretable, voxel-level evidence directly from raw fMRI inputs.
@@ A.1 Preprocessing
HCP and CHCP data were preprocessed using the HCP minimal preprocessing pipeline (glasser2013minimal). ABCD data were processed using the ABCD-HCP pipeline. PIOP1, PIOP2 and ISYB datasets were preprocessed using fMRIPrep (esteban2019fmriprep). For ABIDE and ADHD, we adopted preprocessed datasets provided by the Preprocessed Connectomes Project (PCP) (craddock2013neuro; bellec2017neuro). To ensure compatibility across datasets, we first resampled all images to a uniform spatial resolution of 2 mm isotropic using cubic B-spline interpolation. For datasets with lower temporal resolution, we further interpolated the time series to a uniform sampling rate of 0.72 s TR, also using cubic B-spline interpolation. Notably, HCP, CHCP, and ABCD natively share nearly identical temporal resolution and already meet the 2 mm spatial resolution requirement after preprocessing.
@@ A.2 Dataset and Downstream Tasks
We utilize the HCP dataset as our primary benchmark for pre-training and initial evaluation, comprising 1,011 subjects. This cohort represents a young adult population with ages ranging from 22 to over 36 years. Specifically, the age distribution includes distinct groups: 22–25 ($N=218$), 26–30 ($N=437$), 31–35 ($N=346$), and 36+ ($N=10$). The dataset maintains a balanced sex ratio with 469 males and 542 females, providing high-quality, standardized data ideal for establishing baseline model performance. We only use one session per participants van2013wu. HCP Fingerprint Fingerprint serves as a subject-identification benchmark. We randomly chose 100 subjects as distinct classes. Three different sessions from the same subjects are assigned to train, validation, and test splits, ensuring that the task evaluates cross-session fingerprint consistency rather than session-specific artifacts.
@@
To evaluate model generalization on a larger scale and a distinct demographic, we incorporate the ABCD dataset, which includes 2,398 subjects casey2018adolescent. In contrast to HCP, this cohort focuses on early brain development, with a tightly clustered mean age of $9.925\pm 0.627$ years. And the dataset features a highly balanced sex distribution (1,226 males and 1,172 females).
@@
ADHD-200 is an open multi-site neuroimaging resource with resting-state fMRI (rs-fMRI) across multiple contributing centers adhd2012adhd. For this task we merge all ADHD subtypes into a single ADHD label and it contains 938 participants (ADHD 356 vs. control 582).
@@
The Autism Brain Imaging Data Exchange is a multi-site, open-access initiative that aggregates structural and resting-state fMRI scans - alongside rich phenotypic data - from individuals with autism spectrum disorder (ASD) and matched typically developing controls to accelerate reproducible neuroimaging research di2014autism. It contains 871 participants (mean age 16.9, std 7.5). We perform regression on age and scale both the labels and results to the range [0, 1] using StandardScaler. Additionally, we divide the age variable into four balanced groups for the classification task.
@@
The Alzheimer’s Disease Neuroimaging Initiative is a large‐scale, longitudinal, multi‐center project initiated in 2004, designed to collect and publicly share clinical data from cognitively normal older adults, individuals with mild cognitive impairment (MCI) and patients with Alzheimer’s disease (AD) jack2008alzheimer. For our task, we separated disorders into two classification with healthy controls.
@@
The Parkinson’s Progression Markers Initiative is an ongoing, international, longitudinal observational study launched in 2010 by The Michael J. Fox Foundation to identify and validate biomarkers of Parkinson’s disease (PD) risk, onset, and progression. We combined two different diseases and healthy controls from the PPMI dataset to create a three-class classification task marek2011parkinson. This cohort consists of 474 subjects (272 males, 202 females; mean age 66.44). We utilize the data for a three-class disease classification task: Prodromal ($303$), PD ($142$), and Control ($29$).

### v2

@@ 1 Introduction
More recently, there has been a growing interest in developing deep learning based foundation models for fMRI data analysis, inspired by their remarkable performance in Computer Vision and Natural Language Processing tasks. The idea is to pre-train (usually in a self-supervised way) a large model to learn general-purpose representations of brain activity on vast neuroimaging data, which can then be adapted to diverse downstream tasks with limited labeled data (e.g., via fine-tuning). Broadly speaking, current efforts on building foundation models, or more generally applying deep learning techniques in fMRI data analysis can be categorized into two threads: atlas-based and atlas-free approaches. The former continues the traditional paradigm of summarizing voxel-level signals into predefined brain regions (Caro et al., 2023; Assran et al., 2023; Yang et al., 2024), therefore leveraging anatomical priors and improving interpretability. However those methods bear several key limitations: i) there is no universally optimal atlas: performance on different downstream tasks may highly depend on parcellation choices, where results across studies using heterogeneous atlas are not directly comparable (Wang et al., 2025; Salehi et al., 2020); ii) averaging based on any pre-defined atlas will inevitably discard important voxel-level information - as a result such models often need to be trained on very large cohorts (e.g., $\sim 60k$ sessions) to perform well (Dong et al., 2024; Caro et al., 2023) (see Fig. 1); and iii) any analysis using such models will be confined to the parcel resolution of the atlas used, where probing within-region structures (e.g., isolating amygdalar subnuclei during fear conditioning) is impossible (Wen et al., 2022).
@@
On the other hand, atlas-free methods (Zhao et al., 2018; Vu et al., 2020; Nguyen et al., 2020) aim to learn directly from voxel-level data without imposing ROI boundaries of certain atlas, allowing the learned models to capture fine-grained functional patterns and potentially discover novel brain organization. However, existing atlas-free approaches have typically been developed for specific tasks rather than as general-purpose foundation models, due to their prohibitively high training cost. The deep learning architectures underlying those models, such as the widely adopted Vision Transformers (ViT), incur quadratic cost (both memory and compute) with respect to input dimensions. This makes it nearly impossible to pre-train them on large-scale fMRI data, and thus building atlas-free foundation models still remains an open challenge. Recent work has explored efficient variants such as Shifted-window (Swin) Transformer (Kim et al., 2023; Sun et al., 2025; Kwon et al., 2024; Peng et al., 2025b), but at each timestamp they still feed the entire dense fMRI volume into the encoder, wasting resources on those voxels ($\approx$70%) outside the brain with no valid signal. Similar inefficiencies also arise along the temporal axis, where existing approaches either only train with data pertain to specific task states, e.g., extracting $\sim$30 timestamps (Shi et al., 2023; Sun et al., 2025; Rosenman et al., 2024), or consider sliding windows with limited sizes, where data within each window is processed independently through the model, with results aggregated thereafter (Kim et al., 2023)(e.g., as shown in Fig. 2b). In practice neither is ideal: the former obviously hurts generalization capabilities of the model beyond task states, while the latter evenly processes every window - may lose focus on those truly important data segments - leading to inferior performance comparing to some of the recent atlas-based approaches in out-of-distribution cases (Dong et al., 2024).
@@ 2 Related Work
Foundation Models for fMRI Analysis. fMRI signals reflect ongoing brain states and cognitive processes. Early work framed decoding as supervised classification or regression on activity patterns, which often produced task-specific features with limited out-of-distribution generalization (Song et al., 2008; Horikawa and Kamitani, 2017; Ye et al., 2023; Zhao et al., 2018; Vu et al., 2020; Nguyen et al., 2020). More recently, the field has pivoted to task-agnostic foundation models trained with self-supervised pipelines on unlabeled data, aiming for representations that transfer across tasks and datasets. One prominent line of work pools voxel signals within a predefined atlas to obtain region-wise time series and then learns objectives on those ROI tokens—either generative (masked reconstruction, as in BrainLM) or latent prediction (as in Brain-JEPA) (Caro et al., 2023; Dong et al., 2024). These approaches are memory-efficient and scalable, but their quality can depend on the chosen parcellation and its granularity. A related direction converts the atlas-reduced time series into pairwise or higher-order functional graphs and optimizes unsupervised objectives on these structures (Yang et al., 2024; Thapaliya et al., 2024; Han et al., 2025). For example, BrainMass augments networks by randomly dropping time points from the BOLD signal during training to encourage robustness (Yang et al., 2024), and hypergraph formulations have been explored to capture higher-order relationships (Han et al., 2025). Complementing these atlas-based routes, volumetric encoders operate directly on 4-D fMRI volumes, thereby avoiding atlas-induced bias and preserving fine spatial detail (Peng et al., 2025a; Sun et al., 2025; Kwon et al., 2024; Kim et al., 2023).In practice, computational and memory constraints often necessitate training on short, task-aligned excerpts rather than full-length recordings. Typical applications include task classification (Shi et al., 2023), brain-state decoding (Sun et al., 2025), and stress prediction (Rosenman et al., 2024). Turning these systems into a general-purpose foundation model remains a major challenge.
@@
Efficient Vision Transformers. Vanilla Vision Transformers (ViTs) apply global self-attention over all tokens. The cost scales quadratically with token count and the model offers weak spatial inductive bias (Dosovitskiy et al., 2020). This is impractical for high-resolution or long 4-D fMRI sequences, where a single scan can yield millions of voxel tokens across time. Hierarchical Transformers address these issues by progressively downsampling tokens and widening channels (e.g., MViT, Swin) (Fan et al., 2021; Liu et al., 2021). For example, Swin restricts attention to local windows and cyclically shifts the partition between layers, which stabilizes memory and accuracy on dense inputs. However, it limits global context to multi-layer message passing, ties the model to a dense, regular lattice, and introduces engineering overhead that complicates masked-token pretraining. These drawbacks are magnified in fMRI: spatial grids are high resolution and a large fraction of voxels are non-brain background. Hiera shows that many hand-crafted components are unnecessary. A minimal hierarchical ViT by pretext task achieves superior speed/accuracy trade-offs and is naturally compatible with MAE-style pretraining (Ryali et al., 2023). For voxel-level 4-D fMRI, this is particularly attractive: with MAE-style masking, non-informative background can be excluded from the encoder, reducing GPU memory and improving throughput while preserving fine-grained brain signals.
@@ 4.1 Dataset
We leverage self-supervised pretraining on four public neuroimaging datasets: a single HCP session (Van Essen et al., 2013), CHCP (Ge et al., 2023), and two releases from the Amsterdam Open MRI Collection (AOMIC)—PIOP1 and PIOP2 (Snoek et al., 2021); the Adolescent Brain Cognitive Development (ABCD) Study Casey et al. (2018). We use 70% of the data (4129 sessions) for training, with the remainder reserved for validation (10%) and testing (20%). External validation spans four datasets: the Autism Brain Imaging Data Exchange (ABIDE) (Di Martino et al., 2014); the ADHD-200 Sample (ADHD) (consortium, 2012); Alzheimer’s Disease Neuroimaging Initiative (ADNI) Jack Jr et al. (2008); Parkinson’s Progression Markers Initiative (PPMI) Marek et al. (2011). See Appendix  A.1 and  A.2 for preprocessing and dataset details.
@@ 4.2 Fine-tuning results
We fine-tune our model and evaluate it on out-of-distribution datasets, as summarized in Table 1 and Appendix A.3. We benchmark against seven representative fMRI models—BrainNetCNN (Kawahara et al., 2017), BrainGNN Li et al. (2021), BrainLM (Caro et al., 2023), BrainMass (Yang et al., 2024), Brain-JEPA (Dong et al., 2024), NeuroSTORM Li et al. (2025) and SwiFT Kim et al. (2023)—using the authors’ released codes or checkpoints (pretrained on 1010k–6464k samples). Note that BrainMass includes these datasets in its pre-training corpus, whereas our method utilizes them strictly for fine-tuning. We adopted the medium-sized model (45M parameters) trained on 4,129 sessions as the default configuration for the experiments. Across all the OOD evaluations, our model consistently outperforms the baselines, indicating that it learns domain-invariant representations that transfer to unseen datasets and tasks with minimal adaptation. Task description (A.2), statistical method (A.3), internal experiment results (A.4), linear probing results results (A.5), configuration details for ours (A.6) and baselines (A.7) are provided in the Appendix.
@@ 4.3 Ablation Study
To avoid ingesting full-length fMRI sequences, we use the global MAE to obtain a coarse descriptor and to score temporal windows; only the top-ranked windows are fed to the 4D encoder (Table 3). We use five consecutive frames to form one window, and multiple such windows are grouped into a set as the input to the voxel-level model. With a per-frame storage layout, we can load arbitrary time indices on demand, cutting $\sim 80\%$ of 4D data I/O and compute when selecting the top $20\%$ of windows, comparing sliding windows average method (Kim et al., 2023). We compare Top-$k$ selection (mutual masked reconstruction) against Random selection under different frame budgets on HCP sex classification with 1K-size pre-trainging model. Top-$k$ consistently achieves higher accuracy and F1 without full-sequence 4D fMRI reads.
@@
We assess the impact of our design decisions on memory and accuracy by 1K-size pretraining model with HCP dataset in Table. 4. First, employing a 4D Hiera encoder with unit-wise attention excludes background and masked units from the forward pass, substantially reducing the memory footprint. Because fMRI signals are not directly interpretable like natural images, we prioritize representation quality over pixel-space fidelity and adopt a JEPA objective, dispensing with a heavy reconstruction decoder. With 200-frame inputs and Top-$k$ selection, our design reduces peak GPU memory from $\sim$8 GB to $\sim$2.4 GB while improving accuracy. Meanwhile, although Hiera–MAE and Hiera–JEPA have comparable per-sample GPU memory footprints, the JEPA variant achieves substantially higher throughput. By contrast, Kim et al. (2023) processes ten 20-frame windows sequentially and averages their predictions; each 20-frame chunk consumes $\approx 3.2\,\mathrm{GB}$ of GPU memory, so handling all windows inflates the per-sample memory footprint. Likewise, Kwon et al. (2024) employ a Swin-UNet–based model for task prediction; although it is not a foundation model and does not reconstruct to the raw signal space, it still requires $\sim 40\,\mathrm{GB}$ of GPU memory to process 30 volumes with a batch size of four. Details are in Appendix A.10
@@ 4.4 Model interpretation
We examine whether SLIM-Brain’s voxel-level representations are neurobiologically meaningful. First, we use Neurosynth—a large-scale, automated meta-analytic platform that aggregates findings from thousands of fMRI studies (Yarkoni et al., 2011)—to obtain disease-associated meta-analytic maps (e.g., ADHD). Next, we derive fine-grained attention maps from the 4D Hiera encoder by projecting token-level attention back to voxel space. (Implementation details are provided in the Appendix A.11.) Integrated-gradients–derived key regions from SLIM-Brain substantially overlap with Neurosynth meta-analytic distributions (Fig. 5). The highlighted regions include dorsolateral prefrontal cortex and anterior cingulate (executive control), inferior parietal and precuneus/posterior cingulate (default mode network), as well as striatal areas, consistent with ADHD-related fronto-striatal and fronto-parietal dysfunctions. These observations indicate that the learned features are not only predictive but also align with established neurobiological patterns, offering interpretable, voxel-level evidence directly from raw fMRI inputs.
@@ A.1 Preprocessing
HCP and CHCP data were preprocessed using the HCP minimal preprocessing pipeline (Glasser et al., 2013). ABCD data were processed using the ABCD-HCP pipeline. PIOP1, PIOP2 and ISYB datasets were preprocessed using fMRIPrep (Esteban et al., 2019). For ABIDE and ADHD, we adopted preprocessed datasets provided by the Preprocessed Connectomes Project (PCP) (Craddock et al., 2013; Bellec et al., 2017). To ensure compatibility across datasets, we first resampled all images to a uniform spatial resolution of 2 mm isotropic using cubic B-spline interpolation. For datasets with lower temporal resolution, we further interpolated the time series to a uniform sampling rate of 0.72 s TR, also using cubic B-spline interpolation. Notably, HCP, CHCP, and ABCD natively share nearly identical temporal resolution and already meet the 2 mm spatial resolution requirement after preprocessing.
@@ A.2 Dataset and Downstream Tasks
We utilize the HCP dataset as our primary benchmark for pre-training and initial evaluation, comprising 1,011 subjects. This cohort represents a young adult population with ages ranging from 22 to over 36 years. Specifically, the age distribution includes distinct groups: 22–25 ($N=218$), 26–30 ($N=437$), 31–35 ($N=346$), and 36+ ($N=10$). The dataset maintains a balanced sex ratio with 469 males and 542 females, providing high-quality, standardized data ideal for establishing baseline model performance. We only use one session per participants Van Essen et al. (2013). HCP Fingerprint Fingerprint serves as a subject-identification benchmark. We randomly chose 100 subjects as distinct classes. Three different sessions from the same subjects are assigned to train, validation, and test splits, ensuring that the task evaluates cross-session fingerprint consistency rather than session-specific artifacts.
@@
To evaluate model generalization on a larger scale and a distinct demographic, we incorporate the ABCD dataset, which includes 2,398 subjects Casey et al. (2018). In contrast to HCP, this cohort focuses on early brain development, with a tightly clustered mean age of $9.925\pm 0.627$ years. And the dataset features a highly balanced sex distribution (1,226 males and 1,172 females).
@@
ADHD-200 is an open multi-site neuroimaging resource with resting-state fMRI (rs-fMRI) across multiple contributing centers consortium (2012). For this task we merge all ADHD subtypes into a single ADHD label and it contains 938 participants (ADHD 356 vs. control 582).
@@
The Autism Brain Imaging Data Exchange is a multi-site, open-access initiative that aggregates structural and resting-state fMRI scans - alongside rich phenotypic data - from individuals with autism spectrum disorder (ASD) and matched typically developing controls to accelerate reproducible neuroimaging research Di Martino et al. (2014). It contains 871 participants (mean age 16.9, std 7.5). We perform regression on age and scale both the labels and results to the range [0, 1] using StandardScaler. Additionally, we divide the age variable into four balanced groups for the classification task.
@@
The Alzheimer’s Disease Neuroimaging Initiative is a large‐scale, longitudinal, multi‐center project initiated in 2004, designed to collect and publicly share clinical data from cognitively normal older adults, individuals with mild cognitive impairment (MCI) and patients with Alzheimer’s disease (AD) Jack Jr et al. (2008). For our task, we separated disorders into two classification with healthy controls.
@@
The Parkinson’s Progression Markers Initiative is an ongoing, international, longitudinal observational study launched in 2010 by The Michael J. Fox Foundation to identify and validate biomarkers of Parkinson’s disease (PD) risk, onset, and progression. We combined two different diseases and healthy controls from the PPMI dataset to create a three-class classification task Marek et al. (2011). This cohort consists of 474 subjects (272 males, 202 females; mean age 66.44). We utilize the data for a three-class disease classification task: Prodromal ($303$), PD ($142$), and Control ($29$).